Bond James Bond Corporation has 7 percent coupon bonds that have 9 years left to maturity. The bonds make annual payments to their holders. The YTM on these bonds is 10 percent. Calculate the current price for each of the Bond James Bond Corporation's bonds.


rev: 09_18_2012

Multiple Choice

  • $1,199.46

  • $868.59

  • $827.23

  • $1,030.00

  • $785.87
Answer #1

Annual coupon=1000*7%=70

Hence current price=Annual coupon*Present value of annuity factor(10%,9)+$1000*Present value of discounting factor(10%,9)

=70*5.759023816+$1000*0.424097618

=$827.23(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=70[1-(1.1)^-9]/0.1

=70*5.759023816

2.Present value of discounting factor=1000/1.1^9

=1000*0.424097618